OpenStudy (anonymous):

Three pounds of butter at $n per pound cost $3.85

OpenStudy (anonymous):

3.85 / 3 = n / 1 Which is the same as: 3.85 / 3 = n 1.28333 = n Rounding 1.28 = n So each pound of butter costs about $1.28.